Solve the differential equation x^3 p^2 +x^2 yp+a^3=0 and also obtain its singular solution, if
it exists.
1
Expert's answer
2020-02-17T10:08:43-0500

"x^3 p^2 +x^2 yp+a^3=0" ---(1) is the given differential equation, where "p=dy\/dx"

"\implies y=-xp-a^3/(x^2p)"

Differentiating with respect to x, we get;

"p=-p-xdp\/dx+a^3\/x^2p^2(dp\/dx)+2a^3\/x^3p"

"\implies 2p+xdp/dx=2a^3/px^3+a^3/x^3p^2(dp/dx)"

"\implies 2p(1-a^3/x^3p^2)+xdp/dx(1-a^3/x^3p^2)=0"

"\implies 2p+xdp/dx=0"

"\implies\int dp/p=-2\int dx/x"

"\implies px^2=c \implies p=c/x^2" ---(2)

Substituting (2) in (1), we get;

"c^2+xyc+a^3x=0" which is a quadratic equation in c;

Equating it's discriminant to 0, we get;

"x^2y^2-4a^3x=0"

"\implies x(xy^2-4a^3)=0"

"\implies x=0 andxy^2-4a^3=0"

Both these satisfy (1).

Thus, "x=0" and "xy^2-4a^3=0" are the singular solutions of the given differential equation.